A pair of fair dice is rolled at a time.the random variable X is the maximum of the two numbers shown on the dice.the mean of X is
Answer by oscargut(2103)

P(X=1) = 1/36
P(X=2) = 3/36
P(X=3) = 5/36
P(X=4) = 7/36
P(X=5) = 9/36
P(X=6) = 11/36
E(X) = 1(1/36)+2(3/36)+3(5/36)+4(7/36)+5(9/36)+6(11/36)
= (1+6+15+28+45+66)/36 = 161/36
Answer: 161/36
I can help you with more questions at: mthman@gmail.com
Thanks